# Linear Discriminant Regularized Regression

### v1

@@ A.2 Proof of Theorem 2
$\mathcal{E}_{\lambda}:=\left\{{1\over n}\|{\mathbb{X}}^{\top}({\mathbb{Y}}-{\mathbb{X}}B)\|_{\infty}\leq{1\over 2}\lambda\right\}$ (A.3)
@@
$\mathcal{E}_{RE}:=\left\{\min_{S\subseteq[p],|S|\leq s}\min_{v\in\mathcal{C}(S,3)}{\|{\mathbb{X}}v\|_{2}^{2}\over n\|v\|_{2}^{2}}\geq{\kappa_{s}\over 2}\right\}.$ (A.4)
@@
We proceed to verify that $\mathbb{P}(\mathcal{E}_{RE})\geq 1-(n\vee p)^{-2}$. To this end, for any set $S\subseteq[p]$, we write $E_{S}:=\text{span}\{{\bm{e}}_{j}:j\in S\}$. Under 2, we define the space
$\Gamma_{s}:=\bigcup_{S\subseteq[p]:\ |S|=\lfloor Cs\rfloor}E_{S}$ (A.5)
for some constant $C=C(\kappa_{s},\|\Sigma_{w}\|_{\infty})>1$. Let $\widehat{\Sigma}=n^{-1}{\mathbb{X}}^{\top}{\mathbb{X}}$. Lemma 7 and the condition $(s\vee L)\log(n\vee p)\leq cn$ imply that
${9\over 10}\|\Sigma^{1/2}v\|_{2}\leq\|\widehat{\Sigma}^{1/2}v\|_{2}\leq{11\over 10}\|\Sigma^{1/2}v\|_{2}\qquad\forall~{}v\in\Gamma_{s}$ (A.6)
holds with probability at least $1-(n\vee p)^{-2}$. According to Rudelson and Zhou (2012, Theorem 3), (A.6) implies that there exists $S\subseteq[p]$ with $|S|\leq s$ such that
$u^{\top}\widehat{\Sigma}u\geq{1\over 2}u^{\top}\Sigma u$ (A.7)
holds for any $u\in\mathcal{C}(S,3)$. Eq. A.7 in conjunction with 2 ensures that $\mathcal{E}_{RE}$ holds.
It is now relatively straightforward to finish the proof. Standard arguments (Bühlmann and Van de Geer, 2011; Giraud, 2021) yield that on the event $\mathcal{E}_{\lambda}$, for any $\ell\in[L]$,
${1\over 2n}\|{\mathbb{X}}\widehat{B}_{\ell}-{\mathbb{X}}B_{\ell}\|_{2}^{2}$ $\leq{\lambda\over 2}\|\widehat{B}_{\ell}-B_{\ell}\|_{1}+\lambda\|B_{\ell}\|_{1}-\lambda\|\widehat{B}_{\ell}\|_{1}$
@@
$\widehat{B}_{\ell}-B_{\ell}\in\mathcal{C}(S_{\ell},3)$ (A.8)
with $S_{\ell}:={\rm supp}(B_{\ell})$. Invoking (A.7), we find on the event $\mathcal{E}_{RE}$
${\kappa_{s}\over 4}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}^{2}\leq{1\over 4}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}^{2}$ $\leq{1\over 2n}\|{\mathbb{X}}\widehat{B}_{\ell}-{\mathbb{X}}B_{\ell}\|_{2}^{2}$
@@
A.2.1 Two key lemmas used in the proof of Theorem 2

### v2

@@ B.4 Proof of 4
$\lesssim\mathbb{E}_{{\bm{D}}}\left[\min\left\{\Delta_{\infty}{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*})+{L\over n},~{}1\right\}1\{\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\}\right]+{1\over n}$
$\lesssim\mathbb{E}_{{\bm{D}}}\left[\min\left\{\Delta_{\infty}{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*}),~{}1\right\}\right],$
completing the proof. ∎
B.5 Proof of Theorem 5
The proof of Theorem 5 uses the following lemma on $\|\widehat{H}-H\|_{\rm op}$. The proof of Lemma 10 appears at the end of this section.
Lemma 10.
Under the conditions of Theorem 5, with probability at least $1-n^{-1}$,
$\|\widehat{H}-H\|_{\rm op}\lesssim\sqrt{\log(n)\over nL}+{\omega_{2}\over\sqrt{L}}.$
Proof of Theorem 5.
Define
$\delta=C\sqrt{\log(n)\over nL}+C{\omega_{2}\over\sqrt{L}}$ (B.10)
for some constant $C>0$. We work on the event
$\mathcal{E}_{H}:=\left\{\|\widehat{H}-H\|_{\rm op}\leq\delta\right\}.$
Lemma 10 ensures that $\mathcal{E}_{H}$ holds with probability at least $1-n^{-1}$. Moreover, on the event $\mathcal{E}_{H}$, we find that
$=\left[\lambda_{1}(\Omega)\right]^{-1}-\|\widehat{H}-H\|_{\rm op}$ by $\text{by }\Omega=H^{-1}$
[…]
Hence, on the event $\mathcal{E}_{H}$, $\widehat{H}$ is invertible. We proceed to prove
$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}^{2}~{}\lesssim~{}\left(\omega_{2}^{2}+\delta^{2}L\Delta_{\infty}\right)L^{2}\Delta_{\infty}^{2},$ (B.11)
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}\lesssim\left(\omega_{1}^{2}{\log(n\vee p)\over n}+\delta^{2}L\Delta_{\infty}{\log(n)\over n}\right)L^{3}\Delta_{\infty}^{2}$ (B.12)
which, in conjunction with (B.10), $\Delta\asymp 1$, $L\log(n)\leq cn$ and $\omega_{2}\sqrt{L}\leq c$, yield the claim.
To prove (B.11), we write $\widehat{\Omega}=\widehat{H}^{-1}$ for simplicity. By definition,
$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H\widehat{\Omega}-\mathbf{I}_{L})$ by $\text{by }H\Omega=\mathbf{I}_{L}$
$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H-\widehat{H})\widehat{\Omega}$ by $\text{by }\widehat{H}\widehat{\Omega}=\mathbf{I}_{L}.$ (B.13)
It then follows that
$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}$
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\widehat{\Omega}\|_{F}+\|\Sigma_{w}^{1/2}B^{*}(H-\widehat{H})\widehat{\Omega}\|_{F}$
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}\|\widehat{\Omega}\|_{\rm op}+\|\Sigma_{w}^{1/2}B^{*}\|_{F}~{}\|(H-\widehat{H})\widehat{\Omega}\|_{\rm op}$
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}\|\widehat{\Omega}\|_{\rm op}+\sqrt{L\Delta_{\infty}}\|\widehat{H}-H\|_{{\rm op}}\|\widehat{\Omega}\|_{\rm op}.$
In the last step, we use $\|\Sigma_{w}^{1/2}B^{*}\|_{F}^{2}={\rm tr}(B^{*\top}\Sigma_{w}B^{*})={\rm tr}({M}^{\top}\Sigma_{w}^{-1}{M})\leq L\Delta_{\infty}$. We bound $\|\widehat{\Omega}\|_{\rm op}$ as follows:
$\|\widehat{\Omega}\|_{\rm op}$ $\leq\|\Omega\|_{\rm op}+\|\widehat{\Omega}-\Omega\|_{\rm op}$
$=\|\Omega\|_{\rm op}+\|\Omega(H-\widehat{H})\widehat{\Omega}\|_{\rm op}$
$\leq\|\Omega\|_{\rm op}+\|\Omega\|_{\rm op}\|H-\widehat{H}\|_{\rm op}\|\widehat{\Omega}\|_{\rm op}.$
Since the event $\mathcal{E}_{H}$ and $\omega_{2}\sqrt{L}\leq c$ ensure that
$\|\Omega\|_{\rm op}\|H-\widehat{H}\|_{\rm op}\leq\left(\max_{k\in[K]}{1\over\pi_{k}}+\Delta_{\rm op}\right)\delta\lesssim\delta L\Delta_{\infty}\leq{1\over 2},$
we conclude
$\|\widehat{\Omega}\|_{\rm op}\leq 2\|\Omega\|_{\rm op}\lesssim L\Delta_{\infty}.$ (B.14)
It now follows from 2 that
$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}$ $\lesssim\left(\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}+\delta\sqrt{L\Delta_{\infty}}\right)L\Delta_{\infty}$
$\leq\left(\omega_{2}+\delta\sqrt{L\Delta_{\infty}}\right)L\Delta_{\infty},$ (B.15)
proving (B.11).
We now prove (B.12). Using the identity (B.5), we have, on the event $\mathcal{E}_{H}$,
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}$
$\leq 2\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}\right]^{2}+2\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}.$ (B.16)
For the first term on the right, we find
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}\right]^{2}$ $=\sum_{\ell=1}^{L}(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}\widehat{\Omega}_{\ell}^{\top}(\widehat{B}-B)^{\top}(\widehat{\mu}_{\ell}-\mu_{\ell})$
$\leq\max_{\ell\in[L]}\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}^{2}\sum_{\ell=1}^{L}\|(\widehat{B}-B)\widehat{\Omega}_{\ell}\widehat{\Omega}_{\ell}^{\top}(\widehat{B}-B)^{\top}\|_{1}.$
On the one hand, on the event $\mathcal{E}_{\pi}$, Lemma 22 with $k=\ell$ and $v={\bm{e}}_{j}$, in conjunction with union bounds over $j\in[p]$ and $\ell\in[L]$, yields that
$\max_{\ell\in[L]}\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}^{2}\lesssim{\|\Sigma_{w}\|_{\infty}}{L\log(n\vee p)\over n}$ (B.17)
with probability at least $1-n^{-2}$. On the other hand, we notice that
$\sum_{\ell=1}^{L}\|(\widehat{B}-B)\widehat{\Omega}_{\ell}\widehat{\Omega}_{\ell}^{\top}(\widehat{B}-B)^{\top}\|_{1}$ $=\sum_{\ell=1}^{L}\sum_{i,j\in[p]}\left|{\bm{e}}_{i}^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}\widehat{\Omega}_{\ell}^{\top}(\widehat{B}-B)^{\top}{\bm{e}}_{j}\right|$
$\leq\sum_{i,j\in[p]}\left(\sum_{\ell=1}^{L}({\bm{e}}_{i}^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell})^{2}\right)^{1/2}\left(\sum_{\ell=1}^{L}({\bm{e}}_{j}^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell})^{2}\right)^{1/2}$
$=\sum_{i,j\in[p]}\|\widehat{\Omega}(\widehat{B}-B)^{\top}{\bm{e}}_{i}\|_{2}\|\widehat{\Omega}(\widehat{B}-B)^{\top}{\bm{e}}_{j}\|_{2}$
$\leq\|\widehat{\Omega}\|_{\rm op}^{2}\sum_{i,j\in[p]}\|(\widehat{B}-B)^{\top}{\bm{e}}_{i}\|_{2}\|(\widehat{B}-B)^{\top}{\bm{e}}_{j}\|_{2}$
$\leq\|\widehat{\Omega}\|_{\rm op}^{2}\|\widehat{B}-B\|_{1,2}^{2}.$
By invoking 2 and applying inequalities (B.14) and (B.17), we obtain
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}\lesssim\omega_{1}^{2}{\|\Sigma_{w}\|_{\infty}}{\log(n\vee p)\over n}L^{3}\Delta_{\infty}^{2}$ (B.18)
with probability at least $1-n^{-2}$. Regarding the second term on the right in (B.5), we have
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}$ $\leq\sum_{\ell=1}^{L}\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{2}^{2}\|(H-\widehat{H})\widehat{\Omega}_{\ell}\|_{2}^{2}$
$\leq\max_{\ell\in[L]}\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{2}^{2}\|H-\widehat{H}\|_{\rm op}^{2}\sum_{\ell=1}^{L}\|\widehat{\Omega}_{\ell}\|_{2}^{2}$
$\leq\delta^{2}\max_{\ell\in[L]}\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{2}^{2}\|\widehat{\Omega}\|_{F}^{2}$
$\lesssim\delta^{2}\max_{\ell\in[L]}\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{2}^{2}~{}L\|\widehat{\Omega}\|_{\rm op}^{2}$
on the event $\mathcal{E}_{H}$. Further invoking (B.7) and (B.14) yields that
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}$ $~{}\lesssim\delta^{2}{\log(n)\over n}L^{4}\Delta_{\infty}^{3}$ (B.19)
with probability at least $1-n^{-2}$. Combining (B.18) with (B.19) proves (B.12), and hence completes the proof. ∎
Proof of Lemma 10.
Recall that
$\begin{split}\widehat{H}-H&=D_{\widehat{\pi}}-D_{\pi}-\left(\widehat{B}^{\top}\widehat{\Sigma}\widehat{B}-B^{\top}\Sigma B\right)\\ &=D_{\widehat{\pi}}-D_{\pi}-B^{\top}(\widehat{\Sigma}-\Sigma)B-(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\\ &\qquad-(\widehat{B}-B)^{\top}\widehat{\Sigma}B-B^{\top}\widehat{\Sigma}(\widehat{B}-B).\end{split}$ (B.20)
By triangle inequality, we have
$\|\widehat{H}-H\|_{\rm op}$ $\leq\|D_{\widehat{\pi}}-D_{\pi}\|_{\rm op}+\|(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\rm op}+2\|(\widehat{B}-B)^{\top}\widehat{\Sigma}B\|_{\rm op}$
$\quad+\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}.$
We proceed to bound each term on the right-hand-side separately. On the event $\mathcal{E}_{\pi}$, we first find that
$\|D_{\widehat{\pi}}-D_{\pi}\|_{\rm op}\leq\max_{\ell\in[L]}|\widehat{\pi}_{\ell}-\pi_{\ell}|\lesssim\sqrt{\log(n)\over nL}.$ (B.21)
Regarding the second term, 2 ensures that
$\|(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\rm op}\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{F}^{2}\leq\omega_{2}^{2}.$ (B.22)
For the last two terms, invoking Lemma 19 ensures with probability at least $1-n^{-2}$,
$\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}\lesssim{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}{\Delta_{\infty}\log(n)\over n}.$ (B.23)
It also follows that
$\|(\widehat{B}-B)^{\top}\widehat{\Sigma}B\|_{\rm op}$ $\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{\rm op}\|\widehat{\Sigma}^{1/2}B\|_{\rm op}$
$\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{\rm op}\left(\|B^{\top}\Sigma B\|_{\rm op}+\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}\right)^{1/2}$
$\lesssim\omega_{2}\sqrt{\Delta_{\rm op}\over L(L+\Delta_{\rm op})}$
where the last step uses
[…]
deduced from Lemma 14. Combining the above bounds with (B.21), (B.22) and (B.23) gives
$\|\widehat{H}-H\|_{\rm op}$ $\lesssim\sqrt{\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}$
$\qquad+{\Delta_{\rm op}\over L+\Delta_{\rm op}}{\Delta_{\infty}\log(n)\over n}+\omega_{2}\sqrt{\Delta_{\rm op}\over L(L+\Delta_{\rm op})}+\omega_{2}^{2}$
with probability at least $1-n^{-1}$. The result follows by using $\Delta_{\infty}\asymp 1$ and $\omega_{2}\sqrt{L}\leq c$ to collect terms. ∎
B.6 Proof of Theorem 6: the lasso estimator of $B$
The proof of Theorem 6 uses the following Restricted Eigenvalue condition (RE) on the within-class covariance matrix $\Sigma_{w}$.
Definition 1.
For any integer $1\leq s\leq p$, define
We prove Theorem 6 under the following weaker condition than 3.
Assumption 5.
Proof of Theorem 6.
We first prove that
$\begin{split}&\max_{\ell\in[L]}{1\over\sqrt{n}}\|{\mathbf{X}}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}\\ &\max_{\ell\in[L]}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}\\ &\max_{\ell\in[L]}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}~{}\lesssim~{}\kappa_{s}^{-1}\lambda\sqrt{s}\end{split}$ (B.24)
hold with probability at least $1-2n^{-1}$. We work on the intersection of the event
$\mathcal{E}_{\lambda}:=\left\{{1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}\leq{1\over 2}\lambda\right\}$ (B.25)
@@ B.6 Proof of Theorem 6: the lasso estimator of $B$
$\mathcal{E}_{RE}:=\left\{\min_{S\subseteq[p],|S|\leq s}\min_{v\in\mathcal{C}(S,3)}{\|{\mathbf{X}}v\|_{2}^{2}\over n\|v\|_{2}^{2}}\geq{\kappa_{s}\over 2}\right\}.$ (B.26)
@@
We proceed to verify that $\mathbb{P}(\mathcal{E}_{RE})\geq 1-(n\vee p)^{-2}$. To this end, for any set $S\subseteq[p]$, we write $E_{S}:=\text{span}\{{\bm{e}}_{j}:j\in S\}$. Under 5, we define the space
$\Gamma_{s}:=\bigcup_{S\subseteq[p]:\ |S|=\lfloor Cs\rfloor}E_{S}$ (B.27)
for some constant $C=C(\kappa_{s},\|\Sigma_{w}\|_{\infty})>1$. Recall that $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$. Lemma 12 and the condition $(s\vee L)\log(n\vee p)\leq c~{}n$ imply that
${9\over 10}\|\Sigma^{1/2}v\|_{2}\leq\|\widehat{\Sigma}^{1/2}v\|_{2}\leq{11\over 10}\|\Sigma^{1/2}v\|_{2}\qquad\forall~{}v\in\Gamma_{s}$ (B.28)
holds with probability at least $1-(n\vee p)^{-2}$. According to Rudelson and Zhou (2012, Theorem 3), (B.28) implies that there exists $S\subseteq[p]$ with $|S|\leq s$ such that
$u^{\top}\widehat{\Sigma}u\geq{1\over 2}u^{\top}\Sigma u$ (B.29)
holds for any $u\in\mathcal{C}(S,3)$. Eq. B.29 in conjunction with 5 ensures that $\mathcal{E}_{RE}$ holds.
It is now relatively straightforward to finish the proof of (B.24). Standard arguments (Bühlmann and Van de Geer, 2011; Giraud, 2021) yield that on the event $\mathcal{E}_{\lambda}$, for any $\ell\in[L]$,
${1\over 2n}\|{\mathbf{X}}\widehat{B}_{\ell}-{\mathbf{X}}B_{\ell}\|_{2}^{2}$ $\leq{\lambda\over 2}\|\widehat{B}_{\ell}-B_{\ell}\|_{1}+\lambda\|B_{\ell}\|_{1}-\lambda\|\widehat{B}_{\ell}\|_{1}$
@@
$\widehat{B}_{\ell}-B_{\ell}\in\mathcal{C}(S_{\ell},3)$ (B.30)
with $S_{\ell}:={\rm supp}(B_{\ell})$. Invoking (B.29), we find on the event $\mathcal{E}_{RE}$
${\kappa_{s}\over 4}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}^{2}\leq{1\over 4}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}^{2}$ $\leq{1\over 2n}\|{\mathbf{X}}\widehat{B}_{\ell}-{\mathbf{X}}B_{\ell}\|_{2}^{2}$
@@
We have proved (B.24).
Finally, the claim follows by noting that
$\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{F}\leq\sqrt{L}\max_{\ell\in[L]}\|\widehat{\Sigma}^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2},$
$\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}\leq\sqrt{L}\max_{\ell\in[L]}\|\Sigma_{w}^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}\leq\sqrt{L}\max_{\ell\in[L]}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2},$
$\|\widehat{B}-B\|_{1,2}=\sum_{j=1}^{p}\|\widehat{B}_{j\cdot}-B_{j\cdot}\|_{2}\leq\sum_{j=1}^{p}\|\widehat{B}_{j\cdot}-B_{j\cdot}\|_{1}=\sum_{\ell=1}^{L}\|\widehat{B}_{\ell}-B_{\ell}\|_{1}.$
@@
B.6.1 Two key lemmas used in the proof of Theorem 6